OpenStudy (anonymous):

Discussion board and post your response to the following questions, being sure to address all parts of the question in your answer. Part 1: When writing linear equations, how do you determine which form of a line to use? Part 2: Choose 1 set of points from the choices below. Then, solve the problem and post your solution, showing your steps. Write an equation in point-slope form for the line that passes through one of the following pairs of point

OpenStudy (anonymous):

please help me

OpenStudy (dumbcow):

point slope form is: \[y - y_1 = m(x-x_1)\] slope-intercept form is: \[y = mx +b\] given 2 points you can use either form to get equation of line

OpenStudy (anonymous):

Omg! thank you

OpenStudy (dumbcow):

haha yw:)